# Semantics-Aligned, Curriculum-Driven, and Reasoning-Enhanced Vulnerability Repair Framework

### v1

@@ 5.1. Research Questions
Table 1. Comparison of SeCuRepair with state-of-the-art baselines on BigVul and PrimeVulAVR.
AVR Approach Base Model Training Strategy CodeBLEU (%)
@@
FAVOR (favor, ) CodeT5 SFT 25.78 11.77
VulMaster (vulmaster, ) CodeT5 SFT 26.33 11.62
GPT-4o (achiam2023gpt, ) GPT-4o NA 25.90 23.41
SeCuRepair Qwen2.5-7B SFT 29.62 25.92
SeCuRepair Qwen2.5-7B Reasoning SFT & RL 35.42 30.79
Results. The results of our automatic evaluation, presented in Table 1, show that SeCuRepair outperforms all baselines on both datasets by a significant margin in terms of CodeBLEU. Specifically, SeCuRepair surpasses the best-performing baseline on BigVul, VulMaster, by 34.52%; and surpasses the best-performing baseline on PrimeVulAVR, GPT-4o, by 31.52%. This result demonstrates the superior ability of SeCuRepair to repair vulnerability and generalize to unseen repositories. A Wilcoxon signed-rank test (woolson2007wilcoxon, ) confirms that all of SeCuRepair’s performance gains over the baselines are statistically significant (p ¡ 0.001), underscoring the effectiveness and robust generalization capability of our framework.
The superiority of our full pipeline becomes clear when compared against a standard SFT baseline (i.e., Qwen2.5-7B+SFT). The two approaches represent fundamentally different training philosophies: the baseline uses a single-stage, brute-force SFT on all available data (6,104 examples), whereas SeCuRepair employs a two-stage process. SeCuRepair’s initial SFT stage is highly selective, using only 484 high-quality distilled examples to establish a strong reasoning-based foundation before its RL stage trains on the full dataset. The final SeCuRepair model decisively outperforms the monolithic SFT approach, achieving a 19.58% higher CodeBLEU score on BigVul and an 18.79% improvement on PrimeVulAVR. This result strongly validates that our two-stage tuning process is far more effective than simply SFT on all available data in a single step.
5.2.2. Human Evaluation.
Experiment Setting. To complement automatic metrics, we conduct a human evaluation to assess the correctness of the generated patches. We compare SeCuRepair against the best-performing open-source (VulMaster) and commercial baseline (GPT-4o). We recruited four participants, all of whom have at least two years of experience in software security and C/C++ programming. From the PrimeVulAVR test set, we randomly sampled 309 examples (calculated for a 95% confidence level with a 5% margin of error). For each sample, participants were shown the vulnerable code alongside the generated patch (presented as a diff). The ground-truth patch was also provided for reference. To prevent bias, the outputs from the different models for each sample were presented in a random order. Given the time-intensive nature of formally verifying a patch’s functional correctness, we adopted a widely used proxy for correctness: participants evaluated the extent to which a generated patch preserved the same semantic functionality as the ground-truth solution. Following existing works (mahbub2023explaining, ; zhang2020retrieval, ; zhang2023learning, ; zhang2020learning, ), each sample is rated by all four participants on a 5-point Likert scale (from 1: not similar at all to 5: exactly the same semantics).
In this work, we aim to support developers with useful repair suggestions rather than replace them with fully automated fixes. Accordingly, we treat scores 1–2 as negative, indicating that the generated patch is of poor quality and unsuitable as a draft for developers. Scores 3–5 are considered positive (workable patch drafts): a score of 3 reflects a patch that captures key logical elements of the oracle patch but is incomplete, while scores 4 and 5 correspond to nearly correct and fully correct patches, respectively. Thus, patches with scores $\geq$3 are regarded as workable, aligning with our goal of providing developers with actionable repair suggestions. Recent work (takerngsaksiri2025human, ) also emphasizes the importance of keeping humans in the loop, with automated tools serving primarily as recommenders. Following this perspective, we evaluate how well SeCuRepair delivers workable patch drafts that serve as strong starting points for developers.
Results. The human evaluation confirms the superiority of our approach. SeCuRepair produces the highest proportion of workable patches: 58.0% compared to 50.0% for GPT-4o and 20.0% for VulMaster. Specifically, SeCuRepair outperforms the best-performing baseline, GPT-4o, by 16% in the number of workable patches generated, demonstrating that our specialized training pipeline is more effective for AVR than using a raw larger model alone. The success of SeCuRepair demonstrates that the combination of a reason-then-edit paradigm and syntactic- and semantic-aware RL is a key factor in guiding language models to generate functionally correct and logically sound repairs.
5.2.3. Case Analysis
Figure 5. An example of SeCuRepair performing semantic repair with syntactic different patch with oracle.
The vulnerable function sampled from PrimeVulAVR (CVE-2021-29513 (cveexample, )), contains a critical NULL pointer dereference vulnerability (CWE-476 (cwe476, )). The code at line 7 directly dereferences the descr->fields pointer in a call to PyDict_Next without first verifying that it is not a null pointer. This could lead to a crash if the function is called with an improperly initialized descriptor.
Answer to RQ1: SeCuRepair significantly and consistently outperforms all baselines on both the BigVul and PrimeVulAVR test sets, with improvements of at least 34.52% and 31.52%, respectively (p ¡ 0.001). Moreover, our human evaluation of patch correctness shows that SeCuRepair generates 16% more workable patches than the best-performing baseline, GPT-4o.
5.3. RQ2. Contribution of Reasoning
To isolate the impact of our reason-then-edit approach, we conduct two targeted ablation studies:
  1. (1)
    SFT with vs. without Reasoning: First, we assess whether training on explicit reasoning traces improves patch quality. We compare two SFT models: one trained on our distilled (vulnerable code, reasoning, ground-truth patch) triples, and a baseline model trained only on (vulnerable code, ground-truth patch) pairs. This experiment directly measures the benefit of incorporating reasoning into the SFT stage.
  2. (2)
    RL with vs. without SFT Initialization: Second, we investigate if the reasoning-enhanced SFT stage provides a better starting point for reinforcement learning. We compare two RL training runs: one initialized from our SFT checkpoint ($\theta_{SFT}$), and another that starts directly from the base model. This experiment evaluates the contribution of the SFT warm-up to the stability and final performance of the RL fine-tuning.
@@ 5.2.3. Case Analysis
Table 2. Repair effectiveness of patch-only vs reasoning+patch supervision on subset of BigVul.
@@
5.3.1. SFT with Reasoning Improves Patch Quality
To measure the impact of training on explicit reasoning, we compare two models: SFT (Patch-Only), trained on ground-truth patches, and SFT (Reasoning+Patch), trained on our distilled examples. To ensure a fair comparison and isolate the effect of the reasoning data, the SFT (Patch-Only) model is trained using the ground-truth patches for the exact same subset of vulnerabilities that are distilled for the SFT (Reasoning+Patch) model.
The results in Table 2 show that while standard SFT improves over the base model, incorporating reasoning provides a distinct advantage. SFT (Reasoning+Patch) outperforms SFT (Patch-Only) by a margin of 6.94% on CodeBLEU. This performance gain indicates that training on explicit reasoning helps the model learn to coordinate semantic edits, moving it beyond simple token-level imitation toward a more robust understanding of the repair task.
5.3.2. SFT Initialization is beneficial for RL
To measure the impact of our SFT stage for RL training, we compare the RL training dynamics of a policy initialized from our SFT (Reasoning+Patch) checkpoint against one starting from the base model. Figure 6 plots the learning curves for both, making the differences in stability and final performance clear.
Figure 6. Training curves of the RL stage between two initialized models: original model v.s. the SFT initialized one. The x-axis is RL training steps; the y-axis is the per-step, semantic-aware reward computed from the generated patches against the oracle. Steeper slopes indicate faster learning; higher plateaus correlate with better endpoint performance.
The SFT-initialized policy demonstrates superior performance across all three key metrics: 1) Faster Convergence: The SFT-initialized policy is significantly more sample-efficient, reaching its best reward by step 30, approximately 33% faster than the non-SFT policy, which requires around 40 steps. 2) Greater Stability: The SFT policy’s reward curve rises smoothly with minimal oscillation. In contrast, the non-SFT curve exhibits a noticeable dip between steps 15–20, indicating less stable learning dynamics. 3) Higher Final Performance: The SFT-initialized policy converges to a higher reward plateau (around 0.38) compared to the non-SFT policy (around 0.36). This higher final reward directly correlates with better downstream repair performance.
Answer to RQ2: Incorporating reasoning improves SFT patch quality by 6.94% on CodeBLEU. Furthermore, this SFT stage provides a crucial warm-start for RL, leading to more stable and more effective policy optimization with faster coverage.
5.4. RQ3. The Contribution of Syntactic- and Semantic-aware RL
Our semantic-aware RL stage has three key design elements: the RL optimization algorithm (GRPO), the syntactic- and semantics-aware reward, and the curriculum learning schedule. To quantify the contribution of each, we conduct a progressive ablation study, comparing the following model variants:
  • A0: Base Model. The original, pre-trained Qwen2.5-7B-Instruct without any fine-tuning.
  • A1: SFT with Reasoning. The base model was fine-tuned on distilled reasoning traces.
  • A2: SFT → Syntactic-aware RL. We start from our A1 checkpoint and then apply RL, but with a reward function based only on the BLEU score. This measures the benefit of RL optimization itself without the semantic component.
  • A3: Syntactic-aware RL → Syntactic- and Semantic-aware RL. We enhance the A2 checkpoint by replacing the BLEU-only reward with our full semantics-aware reward function (BLEU + AST + DFG). This isolates the contribution of the semantic reward signal.
  • A4: Full SeCuRepair Model. Finally, we add the curriculum learning schedule to the A3 checkpoint. This measures the incremental benefit of the curriculum learning schedule.
Results. We show three key insights of the effectiveness of our RL pipeline, as illustrated in Fig 7.
@@ 5.3.2. SFT Initialization is beneficial for RL
Figure 7. Step-wise Performance Improvement of SeCuRepair Across Training Stages.
  • GRPO-style RL optimization boost the performance of SeCuRepair under SFT (from A1 to A2). The RL stage allows the model to learn from its own generated outputs and recover from generating patches far off track, a crucial capability that cannot be learned from imitating a single reference patch via SFT. The RL design achieves a 15.89% improvement over the SFT.
  • Semantic-aware reward improves over a purely syntactic signal (from A2 to A3). Our composite reward aligns the training objective with functional correctness by incorporating structural (AST) and data-flow (DFG) similarity. This guides the model to learn the underlying behavior of a correct patch, not just its lexical form, leading to 7.67% improvement in CodeBLEU.
@@
  • Difficulty-aware curriculum adds performance boost (from A3 to A4). Introducing the curriculum learning schedule improves the performance of SeCuRepair by 2.97%. Although the aggregate gain is modest, a per-bucket analysis reveals that the curriculum yields substantial improvements precisely where they are needed most: functions with 2-10 vulnerable regions by 4.24% and functions with ¿10 vulnerable regions by 9.46% in terms of CodeBLEU.
Answer to RQ3: The ablation study highlights that each stage of SeCuRepair contributes to the final performance. Specifically, GRPO-style RL optimization provides a 15.89% improvement over the SFT adopted in prior works. The semantics-aware reward adds a 7.67% gain, while difficulty-aware curriculum learning enhances the model’s ability to fix multi-hunk vulnerabilities by 9.49% for functions with more than 10 vulnerable regions.
6. Discussions
6.1. Discussion of Most Dangerous CWE Performance

### v2

@@ 6 Results
Table 1: Comparison of SeCuRepair with state-of-the-art baselines on BigVul and PrimeVulAVR.
Approach Strategy CodeBLEU (%)
@@
FAVOR SFT 25.78 11.77
VulMaster SFT 26.33 11.62
GPT-4o NA 25.90 23.41
SeCuRepair SFT 29.62 25.92
SeCuRepair SFT & RL 36.13 31.27
Table 2: Human Preference Distribution across Different AVR Systems.
Score SeCuRepair GPT-4o VulMaster
1 9 28 83
2 121 127 164
3 124 104 47
4 24 31 12
5 31 19 3
AVG 2.83 2.63 1.99
@@
Table 3: Repair effectiveness of patch-only vs reasoning+patch supervision on BigVul.
@@
The human evaluation result in Table 2 confirms the superiority of our approach with moderate Kappa agreement. SeCuRepair achieves the highest rate of workable patches (58.0%), outperforming GPT-4o (50.0%) and VulMaster (20.0%). This result validates that a specialized training pipeline is more effective for AVR than relying solely on large-scale generalist models. Qualitative analysis further reveals that SeCuRepair frequently generates patches that are syntactically distinct yet semantically identical to human references, demonstrating its capacity to learn semantic repair logic rather than token-level mimicry. (see Appendix I.2 for case study). Meanwhile, patches rated as “workable”, despite minor issues like over-defensive logic or verbose syntax, successfully capture the underlying fix semantics (see Appendix I.3 for detailed definition and criteria).
@@
Figure 3: Step-wise ablation performance of SeCuRepair across training stages. We confirm that each technical design contributes to the final performance.
3) Each technical design progressively contributes to the final performance. SeCuRepair trains LLMs in a multi-stage process. Therefore, we conduct a progressive ablation study on each design to quantify their contributions, as shown in Figure 3. Specifically, we compare the following model variants: 1) Base Model (A0). The original, pre-trained Qwen2.5-7B-Instruct without any fine-tuning. 2) SFT with Reasoning (A1). The base model was fine-tuned on distilled reasoning traces. 3) SFT → Syntactic-aware RL (A2). We start from our A1 checkpoint and then apply RL, but with a reward function based only on the BLEU score. 4) Syntactic-aware RL → Syntactic- and Semantic-aware RL (A3). We enhance the A2 checkpoint by replacing the BLEU-only reward with our full semantics-aware reward function (BLEU + AST + DFG + CFG). 5) Full SeCuRepair Model (A4). Finally, we add the curriculum learning schedule to the A3 checkpoint We confirm that each design progressively contributes to the final performance, highlighting the effectiveness of each component and the synergistic nature of SeCuRepair.
4) Reasoning-enhanced initialization is effective. We investigate whether the reasoning-enhanced SFT stage provides a better starting point for reinforcement learning. We compare two RL training runs: one initialized from our SFT checkpoint ($\theta_{SFT}$), and another that starts directly from the base model. Figure 4 plots the learning curves for both settings. We observe SFT-initialized policy outperforms the non-SFT variant by exhibiting faster convergence (reaching peak reward 33% earlier), greater stability (avoiding mid-training volatility), and a higher final reward plateau (0.38 vs. 0.36), which directly correlates with improved downstream repair performance.
Figure 4: Training curves of the RL stage.
5) Semantic-aware reward improves repairing performance over lexical reward. As shown in Figure 3, GRPO with our semantic-aware reward performs better than using a lexical reward signal that compares the string similarity to the human reference based on BLEU(from A2 to A3 checkpoint). This reward design helps the model to learn from syntactically different but semantically similar patches to the human reference, guiding the model to learn the underlying repair logic over lexical mimicry, leading to 7.67% relative improvement in CodeBLEU.
@@
6) Difficulty-aware curriculum adds performance boost. We observe from Fig 3 (from A3 to A4), introducing a curriculum learning schedule improves the performance of SeCuRepair by 2.97%. A per-bucket analysis reveals that the curriculum yields substantial improvements on complex fix: functions with 2-10 code hunks by 4.24% and functions with >10 vulnerable regions by 9.46% in terms of CodeBLEU. In addition, we compare the performance of SeCuRepair and baselines across different count hunk in Figure 5, we observe SeCuRepair consistently outperforms baselines in all hunk counts, highlighting the effectiveness of this curriculum training strategy.
Figure 5: CodeBLEU across different hunk counts.
7 Analysis
1) SFT with vs. without Reasoning. We assess whether training on explicit reasoning traces improves patch quality. We compare two SFT Qwen-2.5-7B: one trained on our distilled (vulnerable code, reasoning, ground-truth patch) triples, and the other model trained on the same dataset without reasoning text, i.e., (vulnerable code, ground-truth patch). The results in Table 3 show that while standard SFT improves over the base model, incorporating reasoning provides a distinct advantage. SFT (Reasoning+Patch) outperforms SFT (Patch-Only) by a margin of 6.94% on CodeBLEU. This performance gain indicates that training on explicit reasoning helps the model learn to coordinate semantic edits, moving it beyond simple token-level imitation toward a more robust understanding of the repair task.
2) Performance on Critical Vulnerabilities. We further analyze performance across the MITRE Top-10 most dangerous CWEs 19. As shown in Table 4, SeCuRepair consistently outperforms the SFT baseline across nearly all categories. The gains are particularly pronounced for high-impact injection vulnerabilities: SQL Injection (CWE-89) improves from 0.201 to 0.430, and XSS (CWE-79) from 0.232 to 0.303. These results confirm that our semantics-aware RL is highly effective at capturing the precise structural patterns required to mitigate complex security flaws. Full results are available at Appendix J.
3) Robustness to Code Perturbation. We perform an ablation study and confirm that SeCuRepair is robust to code perturbation. Following the existing code perturbation strategy Yang et al. (2022), we replace the first local variable in the input with a randomized string to create a syntactically valid variant. Since local variable names do not dictate control flow, this modification guarantees that the code syntax remains strictly equivalent, with minimal affect to code semantics. On the perturbed PrimeVul dataset, VulMaster’s CodeBLEU dropped by 14.75%, while SeCuRepair declined only 3.39%, highlighting the superior robustness of SeCuRepair to code perturbation.
8 Conclusion
We reveal that current AVR approaches struggle with syntactic brittleness and multi-hunk complexity. In response, we propose SeCuRepair, which leverages a reason-then-edit mechanism, semantics-driven RL, and curriculum-training strategy to ground repairs in diagnostic planning and structural correctness. SeCuRepair achieves state-of-the-art performance on BigVul and our proposed PrimeVulAVR, validating the necessity of aligning training objectives with code semantics rather than surface forms. We plan to broaden SeCuRepair to diverse programming languages in future work.
Limitations
We propose SeCuRepair, an automated vulnerability repair framework that synergizes expert-level reasoning with semantics-aware reinforcement learning and curriculum-training. While SeCuRepair demonstrates significant improvements over state-of-the-art baselines, we acknowledge several primary limitations. First, reliance on static proxy metrics. Our training framework utilizes static semantic rewards rather than execution-based feedback. While execution feedback provides ground-truth verification, it is computationally expensive and brittle in practice. We intentionally relax this constraint to ensure scalability across large-scale, real-world datasets (e.g., BigVul) where build environments are often incomplete. Although strictly "noisier" than dynamic testing, our results confirm that these proxies effectively guide the model toward effective repair. Meanwhile, this reward remains grounded in reference comparison. However, we posit that this approach serves as a middle ground to mitigate syntactic overfitting. It enables SeCuRepair to recognize and reward patches that are syntactically diverse yet structurally faithful to the developer’s intent, marking a crucial step toward fully reference-agnostic semantic repair. Second, language specificity. Our current evaluation is restricted to C/C++ vulnerabilities. Although the underlying reason-then-edit paradigm and semantic-aware RL are designed to be language-agnostic, their generalizability to other ecosystems (e.g., Python, Java) or multi-lingual contexts remains to be empirically validated. Third, dependency on closed-source teachers. Our cold-start reasoning data is distilled from proprietary LLMs (e.g., GPT-4o). The black-box nature of these teachers poses challenges for full reproducibility and may introduce distillation bias. To mitigate these issues, we release our distilled datasets and enforce a strict security-expert aligned workflow to filter low-quality rationales, ensuring the student model learns robust patterns. Lastly, our automatic evaluation relies on static metric of CodeBLEU. To mitigate this issue, we complement with human evaluation, quativative and quantitative case studies, validating SeCuRepair’s effectiveness to help developers.
Ethical Consideration
The primary goal of SeCuRepair is to enhance software security by automating the remediation of software defects. We recognize the theoretical risk that such models could be repurposed for malicious vulnerability discovery. However, given the rapid growth of software vulnerabilities outpacing manual repair capacity, we argue that advancing automated defensive tools is an ethical imperative. All data utilized in this study are publicly available, and we strictly adhere to the usage policies of the respective data sources.
Appendix A Appendix
Appendix B Variable Pertubation
To probe whether AVR approaches learn real semantic repair patterns or merely overfit to lexical cues, we test the robustness of baselines (VulMaster Zhou et al. (2024), FAVOR Dong et al. (2025), and VulMaster’s base LLM CodeT5 Wang et al. (2021)) under a simple refactoring: renaming local variables while preserving code semantics. We randomly selected 10 cases where baseline models (trained and evaluated under random splits) correctly generated a patch, then applied consistent variable renamings. For each instance, we manually renamed locally initialized variables to contextually appropriate alternatives (e.g., swapping len for size), ensuring the code remained natural and semantically equivalent, before re-evaluating the models.
Observation. All studied models are highly sensitive to these minor changes, failing to generate functionally equivalent patches in 40–70% of cases, based on manual inspection by the first author. Figure 1 provides an example of this failure. In the original patch (a), VulMaster correctly inserts a safety guard to prevent a buffer over-read. However, after simply renaming the variable plenbytes to nbytes (b), the model fails. It does not generate the equivalent safety check but instead hallucinates an irrelevant function call, leaving the vulnerability unresolved. This case study provides evidence that the model has not learned the underlying semantic rule, i.e., “the return value of the decoding function must be checked before proceeding.” Instead, it learns a superficial correlation between the specific variable name plenbytes and the patch. While in limited scale, the consistent failures strongly motivates the need for an AVR learning framework that prioritizes syntactic and semantic understanding of repair patterns over simple lexical mimicry.
Appendix C Suboptimal on Multi-Hunk Repair
Figure 6: The performance of existing AVR approaches drops as the number of hunks increases.
Setup. Real-world vulnerability fixes frequently span multiple, non-contiguous regions (hunks) that must be edited consistently. Prior work FAVOR Dong et al. (2025)’s case study notes that multi-hunk repair significantly increases AVR complexity and is a common failure mode for sequence-to-sequence models. To quantify this, we evaluate baseline models on our repository-split test set, stratifying the results by the number of hunks in each function.
Appendix D Cross-Repository Generalization Gap
Figure 7: Performance comparison of VulMaster, FAVOR, and CodeT5 under random split and repository-level split. The performance of selected models degrades substantially under repository-level split by up to 91.6% in terms of Exact Match and up to 29.7% in terms of CodeBLEU.
Setup. AVR models must generalize beyond the repositories seen in training. We formalize this through a cross-repository evaluation protocol. We evaluate the baselines on the common dataset used by VulMaster and FAVOR’s dataset: BigVul dataset Fan et al. (2020). We compare their performance under two distinct data splitting strategies: the conventional random split (using the same split as FAVOR) and our strict repository-level split. In the repository-level split, all functions from a given project are confined to a single set (training, validation, or test). To ensure a fair comparison, we fine-tune all models for each strategy. To prevent data contamination, we exclude the pre-trained adaptor modules of VulMaster, as their training relies on a bug-fixing corpus with known project-level overlaps with the BigVul test set.
Appendix E Dataset
. For fair comparison with prior AVR approaches, we adopt the BigVul corpus Fan et al. (2020) as used by VulMaster Zhou et al. (2024) and FAVOR Dong et al. (2025) for the main evaluation. We construct a repository-level split: all functions from a given repository appear in exactly one of the training, validation, and test sets (no repository overlap). We use an 8:1:1 train/val/test ratio and verify that no repository crosses splits.
To rigorously test how well models generalize to entirely unseen projects, we constructed PrimeVulAVR, a new AVR test dataset derived from PrimeVul Ding et al. (2024), a corpus originally created for vulnerability detection in C/C++. Following the data extraction pipeline from VREPAIR Chen et al. (2022), we created vulnerable-patched function pairs from PrimeVul’s CVE-linked commits. To ensure a truly external test set, we then filtered out any function pair whose repository was present in our BigVul training set. This process resulted in a clean, out-of-distribution test set containing 1,554 C/C++ function pairs for evaluating cross-repository generalization.
Appendix F Implementation Details
We implement SeCuRepair with HuggingFace Transformers 26 for SFT training, vllm 28 for inference, and Verl 27 for RL training. Due to excessive training costs for RL, we select one base model, Qwen2.5-7B-Instruct Bai et al. (2023), as proof-of-concept. Qwen2.5 is commonly used in software engineering tasks Yang et al. (2024); Aggarwal et al. (2025); Ke (2025) and relevant vulnerability detection task Weyssow et al. (2025) as the base model. This setting aligns with existing AVR approaches Chen et al. (2022); Dong et al. (2025), which fine-tune one base model.
For a controlled comparison, all open-source models (our baselines and SeCuRepair) are trained on our BigVul training split. During evaluation, we use deterministic decoding by setting the temperature to 0 for all models.
Appendix G Baselines
We evaluate SeCuRepair against the following groups of baselines:
  • Learning-based AVR approaches. We select two state-of-the-art AVR approaches, VulMaster Zhou et al. (2024) and FAVOR Dong et al. (2025). VulMaster applies CWE expert knowledge to guide the repair process and can handle long input sequences. FAVOR augments the input function with CFG and historical patches. We retrain both on our repository-level split using their recommended hyperparameters.
  • Commercial LLMs. We also evaluate against a top-tier commercial model GPT-4o Achiam et al. (2023) to benchmark against the general-purpose state-of-the-art. We prompt GPT-4o with the same instructional wrapper used for SeCuRepair to ensure a fair, direct comparison of repair capabilities.
  • SeCuRepair base model with SFT. We fine-tune the base model of SeCuRepair, Qwen2.5-7B-Instruct Bai et al. (2023), on our repository-level BigVul training split as one baseline. We select the best checkpoint based on the performance of the validation set.
Appendix H The Brittleness of the Exact Match Metric.
The case study in Figure 1 highlights the fundamental brittleness of the Exact Match (EM) metric for evaluating AVR. EM is fundamentally sensitive to syntactic variations that preserve the code’s semantics. For example, a logically equivalent change like rewriting if (a > b) to if (b < a) would still result in an EM score of 0. Indeed, an ideal AVR approach capable of semantic repairing should be able to generate a diverse set of semantically correct and equivalent patches, most of which would be wrongly evaluated by EM as they are lexically different from the single oracle patch. This strictness makes EM an unreliable indicator of a patch’s correctness and ill-suited for evaluating advanced, semantics-aware AVR approaches. In contrast, CodeBLEU considers AST similarity, making it robust to syntactic variations. Therefore, we consider CodeBLEU a more meaningful metric for AVR and use it as our primary measure in the following sections.
Appendix I Human Evaluation.
I.1 Experiment Setting.
To complement automatic metrics, we conduct a human evaluation to assess the correctness of the generated patches. We compare SeCuRepair against the best-performing open-source (VulMaster) and commercial baseline (GPT-4o). We recruited four participants, all of whom have at least two years of experience in software security and C/C++ programming. From the PrimeVulAVR test set, we randomly sampled 309 examples (calculated for a 95% confidence level with a 5% margin of error). For each sample, participants were shown the vulnerable code alongside the generated patch (presented as a diff). The ground-truth patch was also provided for reference. To prevent bias, the outputs from the different models for each sample were presented in a random order. Given the time-intensive nature of formally verifying a patch’s functional correctness, we adopted a widely used proxy for correctness: participants evaluated the extent to which a generated patch preserved the same semantic functionality as the ground-truth solution. Following existing works Mahbub et al. (2023); Zhang et al. (2020a, 2023, b), each sample is rated by all four participants on a 5-point Likert scale (from 1: not similar at all to 5: exactly the same semantics).
In this work, we aim to support developers with useful repair suggestions rather than replace them with fully automated fixes. Accordingly, we treat scores 1–2 as negative, indicating that the generated patch is of poor quality and unsuitable as a draft for developers. Scores 3–5 are considered positive (workable patch drafts): a score of 3 reflects a patch that captures key logical elements of the oracle patch but is incomplete, while scores 4 and 5 correspond to nearly correct and fully correct patches, respectively. Thus, patches with scores $\geq$3 are regarded as workable, aligning with our goal of providing developers with actionable repair suggestions. 1) Score-3: semantically similar but partial fixes, either incomplete or containing superfluous logic. 2) Score-4: high semantic alignment but differing implementation. 3) Score 5: exact logical match.
Recent work Takerngsaksiri et al. (2025) also emphasizes the importance of keeping humans in the loop, with automated tools serving primarily as recommenders. Following this perspective, we evaluate how well SeCuRepair delivers workable patch drafts that serve as strong starting points for developers.
I.2 Case Analysis
Figure 8: An example of SeCuRepair performing semantic repair with syntactic different patch with oracle.
The vulnerable function sampled from PrimeVulAVR (CVE-2021-29513 6), contains a critical NULL pointer dereference vulnerability (CWE-476 7). The code at line 7 directly dereferences the descr->fields pointer in a call to PyDict_Next without first verifying that it is not a null pointer. This could lead to a crash if the function is called with an improperly initialized descriptor.
I.3 Quantitative Analysis
Human annotators are asked to summarize the common failure modes in those workable but imperfect patches (scored 3 and 4). We observe score-3 patches exhibit three common failure modes: over-defensive fixes (e.g., excessive null checks that may alter intended behavior), minor logic gaps or partially applied fixes, and patches entangled with faulty refactoring. Score-4 patches are semantically similar to the Oracle but differ in implementation, often being syntactically verbose, less readable, over-commented, or using alternative APIs that may have minor semantics differences.
Appendix J Discussion of the most dangerous CWE performance.